help me name this property if 4m=-15,then-12m-45
Answer by orca(409)

First we need to solve the equation 4m = -15 for m.
m= -15/4
Next substitute m = -15/4 into the expression -12m-45.
So
-12m-45 = -12(-15/4)-45=45-45=0
**************************************************************
ALTERNATIVE SOLUTION:
-12m - 45
=-3*(4m)-45
=-3(-15)-45 (plug in 4m = -15)
=45-45
=0